# DiSCo: Device-Server Collaborative LLM-Based Text Streaming Services

### v1

@@ 5 Evaluation
Through extensive experimentation with four production-grade LLM services and state-of-the-art open-source models, we demonstrate DiSCo’s exceptional performance. Our rigorous evaluation spanning diverse deployment scenarios reveals that DiSCo delivers remarkable improvements - reducing both mean TTFT (6-78%) and tail TTFT (11-52%) while achieving cost savings of up to 83.6% compared to existing approaches, all while preserving consistent token generation throughput.
@@ 5.1 Evaluation Setup
Our testbed is a server with 4 NVIDIA A40 GPUs, each featuring 48GB memory. We evaluate DiSCo using both commercial LLM traces and on-device deployments. For server-side evaluation, we collect traces from four production services: OpenAI’s GPT-4o-mini (OpenAI, 2024), DeepSeek-V2.5 (DeepSeek, 2024), Cohere’s Command (Cohere, 2024), and Hyperbolic-hosted LLaMA-3-70b-Instruct (Hyperbolic, 2024). For on-device evaluation, we test three representative device-model configurations (Li et al., 2024b): Pixel 7 Pro with Bloom 1.1B (31.32/13.93 tokens/s for prefill/decode), Pixel 7 Pro with Bloom 560M (51.80/20.14 tokens/s), and Xiaomi 14 with Qwen 1.5 0.5B (79.90/21.47 tokens/s). These configurations span different compute-capability trade-offs in mobile environments. For end-to-end cost comparison, we quantify server costs using commercial API token pricing and device costs using FLOPs-based energy consumption. The detailed cost analysis can be found in Appendix E.
@@
We evaluate the system performance using both TTFT and TBT, including their mean and tail values. They are analyzed across varying cost budgets, defined as the ratio of input tokens processed by the constrained endpoint (device or server) to the total input tokens.
For each experiment, we report the mean value over 10 runs.
@@ 5.2 End-to-end Performance
Figure 6 and Table 2 show that DiSCo significantly outperforms baseline methods in both device- and server-constrained settings, showing improvements across mean and tail (P99) TTFT metrics for various services, including GPT, LLaMA, DeepSeek, and Command. In the GPT experiments, DiSCo demonstrates particularly notable tail latency reductions, decreasing P99 TTFT by up to 40% relative to Stochastic dispatching across all device configurations, while mean TTFT is also reduced substantially, with reductions between 20-30% across diverse budget ratios. In the LLaMA setup, we observe a unique trade-off pattern. For budget ratios below 20% when the device is the constrained endpoint, DiSCo exhibits a slightly higher mean TTFT than the baseline. This outcome is intentional, as DiSCo prioritizes tail latency reduction in low-budget scenarios, yielding substantial gains in P99 TTFT—reducing tail latency by up to 50%. This prioritization enables more responsive performance under restrictive budget conditions.
@@ 5.3 Performance Breakdown and Ablation Study
Impact of Prompt Sending Interval.
To evaluate our system under realistic workload patterns, we conduct experiments using stratified sampling based on request frequency from DiffusionDB (Wang et al., 2022). Specifically, we select traces from ten users across different activity levels to capture diverse interaction patterns. We pair these real-world request intervals with prompts randomly drawn from the Alpaca dataset (Taori et al., 2023). The results shown in Figure 5 demonstrate that DiSCo’s performance advantages persist across varying user activity patterns.
@@
We conduct comprehensive experiments using instruction-following tasks on multiple model configurations. We employ three LLM-based judges (GPT-4o, Gemini1.5-pro, and QWen2.5-72b) to assess response quality, and examine two representative migration scenarios: from a smaller to larger model (3B-7B) and vice versa (7B-3B). Figure 8 shows that DiSCo maintains quality scores across different sequence lengths, migration patterns, and judges. Detailed results are presented in Appendix D.
@@
We conducted comprehensive performance evaluations of DiSCo-D and DiSCo-S on a MacBook Pro with M1 processor, using both synthetic datasets and a real-world GPT trace of 1,000 records, across target frequencies from 0 to 1. To generate synthetic data that accurately reflects real-world scenarios, we fitted log-normal distributions to the prompt lengths and TTFT from the real trace by following the mean and standard deviation of the logarithm. As shown in Figure 9, for DiSCo-S, the execution time showed remarkable efficiency: 0.128 ms for the real trace with 1K samples, scaling to just 0.969 ms and 9.082 ms for synthetic datasets of 10K and 100K samples respectively. DiSCo-D, while being more computationally intensive, still maintained practical performance levels: 0.486 ms, 1.741 ms, and 14.856 ms for 1K, 10K, and 100K samples, respectively.
@@ 6 Conclusions
This paper introduces DiSCo, a device-server cooperative scheduler that addresses QoE and cost challenges in LLM serving for real-time conversational applications for end users. DiSCo uses cost-aware scheduling and token-level migration to dynamically optimize Time-To-First-Token (TTFT) and Time-Between-Token (TBT) across device and server endpoints. Our evaluations on real-world traces from platforms like GPT and DeepSeek show that DiSCo significantly improves both TTFT and TBT while reducing costs.
@@ 7 Limitations
Model Coverage.
We focus on scenarios where on-device LLMs achieve sufficient accuracy for target applications. While this covers many common use cases, DiSCo may not be suitable for applications requiring complex reasoning.
@@
For device energy consumption, we use a linear energy model based on FLOPs. Real-world device energy consumption patterns can be more complex, varying with factors such as battery state, temperature, and concurrent workloads.
Scalability Considerations.
Our current implementation and evaluation focus on single-device scenarios. Extending DiSCo to handle multi-device collaborative serving presents additional challenges in terms of coordination overhead and resource allocation that warrant further investigation.
@@ Appendix A Additional Related Work
LLMs generate text responses auto-regressively, producing one token at a time based on preceding tokens. The process consists of two stages that can potentially be executed on different endpoints: (i) Prefill stage: The model processes the input text (prompt), calculates and stores intermediate model states–i.e., the key and value cache (KV cache) of tokens–to generate the first token. A token represents a word or part of a word the model can interpret. Once the first token is generated, it is appended to the end of the prompt and the generation process moves on to the (ii) Decode stage: The model processes the updated prompt (including previously generated tokens) to generate the next token. The decode stage continues until a stopping condition is met (e.g., reaching an end-of-sequence token or the maximum generation length).
@@
Existing work have focused on GPU kernel optimization (Dao et al., 2022; Ye et al., 2025), KV-cache management (Lin et al., 2024a; Liu et al., 2024d; Qin et al., 2024), model parallelism (Shoeybi et al., 2019; Pope et al., 2023; Liu et al., 2023), quantization (Xiao et al., 2023; Lin et al., 2024b; Dettmers et al., 2022), and scheduling (Yu et al., 2022; Kwon et al., 2023; Agrawal et al., 2024; Sun et al., 2025). For example, Orca Yu et al. (2022) introduced continuous batching to improve serving throughput, while vLLM Kwon et al. (2023) developed PagedAttention to reduce LLM memory restraint. Sarathi (Agrawal et al., 2024) implemented chunked prefill to mitigate inter-request interference within batches. Andes (Liu et al., 2024a) addresses QoE for individual requests from the server side but lacks awareness of network jitter and device potential. These server-side advancements complement DiSCo design.
@@
Google’s Gemini Nano Google (2024) and Apple’s Apple Intelligence Gunter et al. (2024) have been integrated to Android OS and iOS devices, respectively. MLC-LLM MLC team (2023) and llama.cpp (Gerganov, 2024) efficiently deploy various LLM on devices. PowerInfer Song et al. (2023) and PowerInfer-2 Xue et al. (2024b) optimize on-device LLM inference speed by leveraging sparsity in model activations. DiSCo acts as a middle layer to schedule and migrate response generation between servers and devices.
@@ Appendix B Cold Start Evaluation
This section presents cold start performance measurements for the Qwen-2.5 model series across different hardware configurations. The experiments were conducted on two platforms: Windows 10 with NVIDIA RTX 3060 12GB and Linux with NVIDIA A40 48GB. A fixed prompt "How to use GitHub?" was used throughout all experiments. We measured two critical metrics: model loading time and TTFT for Qwen-2.5 models ranging from 0.5B to 7B parameters, all using FP16 precision. The experimental setup consisted of 10 measurement runs, with 2 additional warmup runs to ensure measurement stability. It is worth noting that such warmups can potentially mask the true gap between model loading and prompt prefill time due to various optimizations, including OS page cache. To maintain authentic cold start conditions, we explicitly cleared the CUDA cache and performed garbage collection before each run.
@@ Appendix C Prediction-based Model Selection
This section provides a comparative analysis of several TTFT prediction methods. For selecting the endpoint with a lower TTFT for each request, TTFT prediction is imperative. For on-device inference, TTFT prediction is straightforward, as TTFT exhibits a linear relationship with prompt length. Conversely, on-server inference TTFT is characterized by high variability, rendering prediction challenging. Moreover, the prediction method itself must be computationally efficient, as its overhead also contributes to end-to-end TTFT.
Table 5 presents a comparative analysis of four common lightweight time-series-based prediction methods applied to traces collected from three prevalent LLM services. Our correlation analysis (Table 1) revealed no significant correlation between prompt length and TTFT; thus, prompt length is omitted as a feature in these prediction methods. We demonstrate that none of these methods offers sufficient accuracy for TTFT prediction.
@@ D.2 Evaluation Methodology
  • First endpoint’s maximum generation length varied through [0, 4, 16, 64, 256] tokens
  • Four model combinations: 0.5B-7B, 3B-7B, 7B-0.5B, and 7B-3B (prefix and suffix denote the model sizes of first and second endpoints respectively)
@@ D.3.1 Quality Metrics
Our comprehensive evaluation reveals several key findings:
  • Bounded Quality: The combined sequence quality consistently remains bounded between individual model performance levels
  • Translation Performance: ROUGE-1 scores maintain stability between 0.23 and 0.26
  • Instruction Following: Scores show consistent ranges from 4 to 6
@@ Appendix E Experiment Settings for End-to-end Cost
E.1 FLOPs of On-Device LLMs

### v2

@@ 5 Evaluation
Through extensive experimentation with four production-grade LLM services and state-of-the-art open-source models, we demonstrate DiSCo’s exceptional performance. Our rigorous evaluation spanning diverse deployment scenarios reveals that DiSCo delivers remarkable improvements, reducing both mean TTFT (6-78%) and tail TTFT (11-52%) while achieving cost savings of up to 83.6%.
@@ 5.1 Evaluation Setup
Our testbed is a server with 4 NVIDIA A40 GPUs, each with 48GB of memory. We evaluate DiSCo using both commercial LLM traces and on-device deployments. For server-side evaluation, we collect traces from four production services: OpenAI’s GPT-4o-mini (OpenAI, 2024), DeepSeek-V2.5 (DeepSeek, 2024), Cohere’s Command (Cohere, 2024), and Hyperbolic-hosted LLaMA-3-70b-Instruct (Hyperbolic, 2024). For on-device evaluation, we test three representative device-model configurations (Li et al., 2024b): Pixel 7 Pro with Bloom 1.1B (31.32/13.93 tokens/s for prefill/decode), Pixel 7 Pro with Bloom 560M (51.80/20.14 tokens/s), and Xiaomi 14 with Qwen 1.5 0.5B (79.90/21.47 tokens/s). These configurations span different compute-capability trade-offs in mobile environments. For end-to-end cost comparison, we quantify server costs using commercial API token pricing and device costs using FLOPs-based energy consumption. The detailed cost analysis can be found in Appendix E.
@@
We evaluate the system performance using both TTFT and TBT, including their mean and tail values. They are analyzed across varying cost budgets, defined as the ratio of input tokens processed by the constrained endpoint (device or server) to the total input tokens. For each experiment, we report the mean value over 10 runs.
@@ 5.2 End-to-end Performance
Figure 6 and Table 2 show that DiSCo significantly outperforms baseline methods in both device- and server-constrained settings, showing improvements across mean and tail (P99) TTFT metrics for various services, including GPT, LLaMA, DeepSeek, and Command. In the GPT experiments, DiSCo demonstrates particularly notable tail latency reductions, decreasing P99 TTFT by up to 40% relative to Stochastic dispatching across all device configurations, while mean TTFT is also reduced substantially, with reductions between 20-30% across diverse budget ratios. In the LLaMA setup, we observe a unique trade-off pattern. For budget ratios below 20% when the device is the constrained endpoint, DiSCo exhibits a slightly higher mean TTFT than the baseline. This outcome is intentional, as DiSCo prioritizes tail latency reduction in low-budget scenarios, yielding substantial gains in P99 TTFT—reducing tail latency by up to 50%.
@@ 5.3 Performance Breakdown and Ablation Study
Impact of Prompt Arrival Interval.
To evaluate our system under realistic workload patterns, we conduct experiments using stratified sampling based on request arrival rate from DiffusionDB (Wang et al., 2022). Specifically, we select traces from ten users across different activity levels to capture diverse interaction patterns. We pair these real-world request intervals with prompts randomly drawn from the Alpaca dataset (Taori et al., 2023). The results shown in Figure 5 demonstrate that DiSCo’s performance advantages persist across varied user activity patterns.
@@
We conduct comprehensive experiments using instruction-following tasks on multiple model configurations. We employ three LLM-based judges (GPT-4o, Gemini1.5-pro, and QWen2.5-72b) to assess the response quality, and examine two representative migration scenarios: from a smaller to larger model (3B-7B) and vice versa (7B-3B). Figure 8 shows that DiSCo maintains quality scores across different sequence lengths, migration patterns, and judges. Detailed results are presented in Appendix D.
@@
We conducted comprehensive performance evaluations of DiSCo-D and DiSCo-S on a MacBook Pro with M1 processor, using both synthetic datasets and a real-world GPT trace of 1,000 records, across target frequencies from 0 to 1. To generate synthetic data that accurately reflects real-world scenarios, we fitted log-normal distributions to the prompt lengths and TTFT from the real trace by following the mean and standard deviation of the logarithm. As shown in Figure 9, for DiSCo-S, the execution time showed remarkable efficiency: 0.128 ms for the real trace with 1K samples, scaling to just 0.969 ms and 9.082 ms for synthetic datasets of 10K and 100K samples, respectively. DiSCo-D, while being more computationally intensive, still maintained practical performance levels: 0.486 ms, 1.741 ms, and 14.856 ms for 1K, 10K, and 100K samples, respectively.
@@ 6 Conclusions
This paper introduces DiSCo, a device-server cooperative scheduler that addresses QoE and cost challenges in LLM serving for real-time conversational applications for end users. DiSCo uses cost-aware scheduling and token-level migration to dynamically optimize TTFT and TBT across device and server endpoints. Our evaluations on real-world traces from platforms like GPT and DeepSeek show that DiSCo significantly improves both TTFT and TBT while reducing costs.
@@ 7 Limitations
Model Accuracy.
We focus on scenarios where on-device LLMs achieve sufficient accuracy for target applications. While this covers many common use cases, DiSCo may not be suitable for complex reasoning tasks Guo et al. (2025).
Privacy Protection.
While privacy protection is a key consideration in model selection Chen et al. (2024); Liu et al. (2024b), DiSCo currently assumes users are comfortable with both on-device and on-server deployments.
@@
For device energy consumption, we use a linear energy model based on FLOPs. Real-world device energy consumption patterns can be more complicated, varying on factors such as battery state, temperature, and concurrent workloads Hoque et al. (2015).
Scalability.
Our current implementation and evaluation focus on single-device scenarios. Extending DiSCo to handle multi-device collaborative serving presents additional challenges in terms of coordination overhead and resource allocation that warrant further investigation Niu et al. (2025).
@@ Appendix A Additional Related Work
LLMs generate text responses auto-regressively, producing one token at a time based on preceding tokens. The process consists of two stages that can potentially be executed on different endpoints: (i) Prefill stage: The model processes the input text (prompt), calculates and stores intermediate model states–i.e., the key and value cache (KV cache) of tokens–to generate the first token. A token represents a word or part of a word that the model can interpret. Once the first token is generated, it is appended to the end of the prompt and the generation process moves on to the (ii) Decode stage: The model processes the updated prompt (including previously generated tokens) to generate the next token. The decode stage continues until a stopping condition is met (e.g., reaching an end-of-sequence token or the maximum generation length).
@@
Existing works have focused on GPU kernel optimization (Dao et al., 2022; Ye et al., 2025), KV-cache management (Lin et al., 2024a; Liu et al., 2024e; Qin et al., 2024), model parallelism (Shoeybi et al., 2019; Pope et al., 2023; Liu et al., 2023), quantization (Xiao et al., 2023; Lin et al., 2024b; Dettmers et al., 2022), and scheduling (Yu et al., 2022; Kwon et al., 2023; Agrawal et al., 2024; Sun et al., 2025). For example, Orca Yu et al. (2022) introduced continuous batching to improve serving throughput, while vLLM Kwon et al. (2023) developed PagedAttention to reduce LLM memory restraint. Sarathi (Agrawal et al., 2024) implemented chunked prefill to mitigate inter-request interference within batches. Andes (Liu et al., 2024a) addresses QoE for individual requests from the server side but lacks awareness of network jitter and device potential. These server-side advancements complement DiSCo’s design.
@@
Google’s Gemini Nano Google (2024) and Apple’s Apple Intelligence Gunter et al. (2024) have been integrated into Android OS and iOS devices, respectively. MLC-LLM MLC team (2023) and llama.cpp (Gerganov, 2024) efficiently deploy various LLMs on devices. PowerInfer Song et al. (2023) and PowerInfer-2 Xue et al. (2024b) optimize on-device LLM inference by leveraging sparsity in model activations. DiSCo acts as a middle layer to schedule and migrate response generation between servers and devices.
@@ Appendix B Cold Start Evaluation
This section presents cold start performance measurements for the Qwen-2.5 model series across different hardware configurations. The experiments were conducted on two platforms: Windows 10 with NVIDIA RTX 3060 12GB and Linux with NVIDIA A40 48GB. A fixed prompt "How to use GitHub?" was used throughout all experiments. We measured two critical metrics: model loading time and TTFT for Qwen-2.5 models ranging from 0.5B to 7B parameters, all using FP16 precision. The experimental setup consisted of 10 measurement runs, with two additional warmup runs to ensure measurement stability. It is worth noting that such warmups can potentially mask the true gap between model loading and prompt prefill time due to various optimizations, including OS page cache. To maintain authentic cold start conditions, we explicitly cleared the CUDA cache and performed garbage collection before each run.
@@ Appendix C Prediction-based Model Selection
This section provides a comparative analysis of several TTFT prediction methods. For selecting the endpoint with a lower TTFT for each request, TTFT prediction is imperative. For on-device inference, TTFT prediction is straightforward, as TTFT exhibits a linear relationship with prompt length. Conversely, on-server TTFT is characterized by high variability, posing challenges for prediction. Moreover, the prediction method itself must be computationally efficient, as its overhead also contributes to end-to-end TTFT.
Table 5 presents a comparative analysis of four common lightweight time-series-based prediction methods applied to traces collected from three prevalent LLM services. Our correlation analysis (Table 1) reveals no significant correlation between prompt length and TTFT; thus, prompt length is omitted as a feature in these prediction methods. We demonstrate that none of these methods offers sufficient accuracy for TTFT prediction.
@@
To address potential concerns about the robustness of our server-side TTFT prediction under varying workload conditions, we emphasize that our distribution-based prediction model is specifically designed to handle workload bursts and dynamic patterns. The statistical distribution we maintain inherently captures these patterns, allowing the model to automatically adapt to changing workloads based on historical data. Furthermore, our system architecture supports dynamic distribution updates through a lightweight server API, which can efficiently transmit real-time server workload information using just tens or hundreds of numbers. This design ensures that our prediction model remains responsive and accurate even during significant load spikes or ephemeral workloads.
@@ D.2 Evaluation Methodology
  • Text Summarization: We evaluate 100 data items from the CNN/DM Nallapati et al. (2016) using our structured prompt template, with quality assessment performed by DeepSeek-V3-0324.
  • Story Writing: We evaluate 100 data items from the WritingPrompts Fan et al. (2018) using our structured prompt template, with quality assessment performed by DeepSeek-V3-0324.
@@
  • First endpoint’s maximum generation length varied through different (e.g., [0, 4, 16, 64, 256] tokens).
  • Four model combinations: 0.5B-7B, 3B-7B, 7B-0.5B, and 7B-3B (prefix and suffix denote the model sizes of first and second endpoints, respectively)
@@ D.3.1 Quality Metrics
Our comprehensive evaluation reveals that the combined sequences’ quality consistently remains bounded between individual model performance levels.
@@
(a) Summarization.
(b) Story Writing.
Figure 11: Quality evaluation results with Summarization and Story Writing.
@@ Appendix E Experiment Settings for End-to-end Cost
E.1 Theoretical Cost Modeling
Our unified cost model combines heterogeneous costs (server monetary and device energy) through a dynamic exchange rate $\lambda$, which users adjust based on preferences (e.g., battery level, server budget). We formalize this as:
Total $\text{Total Cost}=\underbrace{c_{s}^{p}\cdot l_{s}+c_{s}^{d}\cdot l_{s}}_{\text{Server Cost}}+\lambda\cdot\underbrace{\left(c_{d}^{p}\cdot l_{d}+c_{d}^{d}\cdot l_{d}\right)}_{\text{Device Cost}}$ (8)
where $l_{s},l_{d}$ denote tokens processed on the server and the device, respectively. The optimization goal is to minimize $\mathbb{E}[\text{Total Cost}]$ while satisfying TTFT/TBT constraints. This formulation allows us to:
  • Balance between server monetary costs and device energy consumption
  • Adapt to user preferences through the exchange rate $\lambda$
  • Account for different costs in prefill ($c^{p}$) and decode ($c^{d}$) phases
  • Optimize token allocation between server and device execution
The model’s flexibility enables it to handle both device-constrained scenarios (where energy consumption is the primary bottleneck) and server-constrained scenarios (where API monetary costs dominate).
While we acknowledge that real-world device energy consumption can be influenced by factors such as battery levels, thermal throttling, and concurrency, we use the linear FLOPs-based measure as a generalizable proxy metric for two key reasons: (1) to cover energy consumption across diverse devices (phones, laptops, and edge servers) and models with varying architectures and quantizations that may exhibit different consumption patterns, and (2) to target short conversational tasks where on-device LLMs demonstrate sufficient efficiency and effectiveness. Importantly, since our dispatching and migration algorithms are both length-threshold-based, they can adapt to any complex energy consumption patterns as long as these patterns are predictable and guaranteed to consume more energy when prefilling or decoding additional tokens.
E.2 FLOPs of On-Device LLMs